# Approaches to critical point theory via sequential and parametrized topological complexity

### v2

@@ 3. Navigation functions and sequential subspace complexities
For each $j\in\{1,2,\ldots,k\}$ we further put $\lambda_{j,1}:=1$, choose a nowhere-vanishing continuous vector field $Y_{j}$ on $\mathbb{S}^{n_{j}}$ and normalize it by putting
$X_{j}(x):=\frac{(r-1)\pi}{\|Y_{j}(x)\|}\cdot Y_{j}(x)\qquad\forall x\in\mathbb{S}^{n_{j}},\ j\in\{1,2,\dots,k\}.$
Note that by this normalization, $X_{j}(x_{j})$ is the tangent vector of a geodesic segment $[a,a+\frac{1}{r-1}]\to\mathbb{S}^{n_{j}}$ from $x_{j}$ to $-x_{j}$ for all $a\in\mathbb{R}$, $x_{j}\in\mathbb{S}^{n_{j}}$, $j\in\{1,2,\dots,k\}$. We define a sequential motion planner $s\colon\Sigma\to PM$ as follows:
@@
  • the constant path in $-x_{j}$ if $\lambda_{j,\ell}=\lambda_{j,(\ell+1)}=-1$,
  • the unique length-minimizing geodesic segment from $x_{j}$ to $-x_{j}$ with $\gamma^{\prime}(\frac{\ell-1}{r-1})=X_{j}(x_{j})$ if $\lambda_{j,\ell}=1$ and $\lambda_{j,(\ell+1)}=-1$,
  • the unique length-minimizing geodesic segment from $-x_{j}$ to $x_{j}$ with $\gamma^{\prime}(\frac{\ell-1}{r-1})=X_{j}(-x_{j})$ if $\lambda_{j,\ell}=-1$ and $\lambda_{j,(\ell+1)}=1$.
One checks that the map $s$ we obtain in this way is indeed a continuous sequential motion planner on $\Sigma$. By Lemma 2.5 extends to a continuous sequential motion planner on an open neighborhood of $\Sigma$. Thus,
$\operatorname{\mathsf{TC}}_{r,M}(\Sigma)=1\quad\forall\Sigma\in\mathcal{C}_{4i}(F),\ i\in\{0,1,\dots,k(r-1)\}.$
Inserting this into (3.4) shows that
$\operatorname{\mathsf{TC}}_{r}(\mathbb{S}^{n_{1}}\times\mathbb{S}^{n_{2}}\times\dots\times\mathbb{S}^{n_{k}})\leq k(r-1)+1$
for all odd numbers $n_{1},n_{2},\dots,n_{k}\in\mathbb{N}$.
@@ 4. Sequential topological complexity of sublevel sets
In this section, we want to derive upper bounds on the sequential TC of a sublevel set of a function $f\colon X\to\mathbb{R}$ from the results of section 2. More precisely, we will apply those results to functions $F\colon X^{r}\to\mathbb{R}$ of the form
@@
and will rephrase the upper bounds in terms of $f$. While we have previously studied subspace complexities of the form $\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})$, it turns out that we can find upper bounds for $\operatorname{\mathsf{TC}}_{r}(f^{\lambda})$, i.e. for the sequential TCs of $f^{\lambda}$ and not just for sequential subspace complexities.
@@
    We begin by considering condition ($D_{1}$). Let $x=(x_{1},\dots,x_{r})\in X^{r}$ with $\Phi(x)\neq x$. Then there exists $i\in\{1,2,\dots,r\}$ with $\varphi(x_{i})\neq x_{i}$. Since $(\varphi,f)$ satisfies ($D_{1}$), this yields $f(\varphi(x_{i}))<f(x_{i})$. Since $f$ is a Lyapunov function for $\varphi$, we derive that
    $F(\Phi(x))=\sum_{j=1}^{r}f(\varphi(x_{j}))<\sum_{j=1}^{r}f(x_{j})=F(x).$
    Thus, $(\Phi,F)$ satisfies ($D_{1}$).
    Concerning ($D_{2}$), we let $B\subset X^{r}$ be such that $\inf\{F(x)-F(\Phi(x))\mid x\in B\}=0$ and such that $F$ is bounded on $B$. Let $(x_{n})_{n\in\mathbb{N}}$ be a sequence in $B$ with
@@
    Moreover, since $(F(x_{n}))_{n}$ is bounded and since $f$ is bounded from below, it follows that $(f(x_{j,n}))_{n}$ is bounded for each $j$. By assumption, $(\varphi,f)$ satisfies condition ($D_{2}$). In particular, it follows that $(x_{1,n})_{n}$ has a convergent subsequence $(x_{1,n_{k}})_{n\in\mathbb{N}}$. Likewise, the sequence $(x_{2,n_{k}})_{k}$ has a convergent subsequence $(x_{2,n_{k_{\ell}}})_{\ell}$. Going on like this and passing to further subsequences until we reach the $r$-th component sequence, we eventually obtain a subsequence of $(x_{n})_{n}$, which we simply denote by $(y_{m})_{m\in\mathbb{N}}$ with the property that all of its component sequences converge. Thus, $(y_{m})_{m\in\mathbb{N}}$ is a convergent subsequence of $(x_{n})_{n\in\mathbb{N}}$ whose limit lies in $\overline{B}\cap\mathrm{Fix\;}\Phi_{1}$ by property ($D_{1}$). This shows that $(\Phi,F)$ has property ($D_{2}$).
@@
    By assumption the restriction $\Phi|_{A}\colon A\to A$ is well-defined. Since $(\Phi,F)$ satisfies condition ($D_{1}$) by a), it is evident that $(\Phi|_{A},F|_{A})$ satisfies ($D_{1}$).
    Let $B\subset A$ be a subset, such that $F$ is bounded on $B$ and
@@
    such that $\operatorname{cl}_{A}(B)\cap\mathrm{Fix\;}\Phi=\overline{B}\cap\mathrm{Fix\;}\Phi\neq\emptyset$, where $\operatorname{cl}_{A}(B)$ denotes the closure of $B$ in $A$ and $\overline{B}$ the closure of $B$ in $X$. It follows from this observation that $(\Phi|_{A},F|_{A})$ satisfies $(D_{2})$.
@@
Apparently, since $f$ is bounded from below, it holds for each $\mu\in\mathbb{R}$ that
$F_{\mu}\cap Y^{r}$ =$=\{(x_{1},\dots,x_{r})\in X^{r}\mid f(x_{1})+f(x_{2})+\dots+f(x_{r})=\mu\}\cap Y^{r}$
$=\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,r\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}(f_{\mu_{i}}\cap Y)=\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}f_{\mu_{i}}$
Since $\mathrm{Fix\;}\Phi=(\mathrm{Fix\;}\varphi)^{r}$, we derive that
$F_{\mu}\cap Y^{r}\cap\mathrm{Fix\;}\Phi=\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}(f_{\mu_{i}}\cap\mathrm{Fix\;}\varphi).$
@@ 5. Sequential parametrized subspace complexities as index functions
Throughout this section let $r\in\mathbb{N}$ with $r\geq 2$ and let $p\colon E\to X$ be a fibration with path-connected fibers. We further let $E^{I}_{X}$, $E^{r}_{X}$ and $\Pi_{r}$ be defined aas in the introduction.
@@
    Let $A\subset E^{r}_{X}$. A local section $s\colon A\to E^{I}_{X}$ of $\Pi_{r}$ is called a sequential parametrized motion planner on $A$
@@
    The $r$-th sequential parametrized subspace complexity of $A$, denoted by $\operatorname{\mathsf{TC}}_{r,p}(A)$, is given as the minimal value of $n\in\mathbb{N}$ for which there exist open subsets $U_{1},U_{2},\dots,U_{n}\subset E^{r}_{X}$ with $A\subset\bigcup_{i=1}^{n}U_{i}$, such that there exists a continuous sequential parametrized motion planner $s_{i}\colon U_{i}\to PX$ for each $i\in\{1,2,\dots,n\}$. In case there is no such $n$, we put $\operatorname{\mathsf{TC}}_{r,X}(A):=\infty$
@@
More abstractly, $q_{r}$ is given as the pullback fibration
$q_{r}=\Delta_{r}^{*}p^{r},$
where $\Delta_{r}\colon X\to X^{r}$, $\Delta(x)=(x,x,\dots,x)$ and where $p^{r}\colon E^{r}\to X^{r}$ denotes the $r$-fold cartesian product of $p$ with itself. In particular, $q_{r}$ is itself a fibration.
@@
    $A$ admits a continuous sequential parametrized motion planner.
@@
such that $s(a)\in E^{I}_{X}$ for all $a\in A$. Thus, $s$ is indeed a continuous sequential parametrized motion planner. ∎
In section 2 we have restricted ourselves to considering ANRs when considering sequential subspace complexities, as we could establish the continuity property of an index function under this assumption. Likewise, we need to make a similar assumption on our fibration $p\colon E\to X$ to ensure that sequential parametrized subspace complexities will have the continuity property. For this purpose, we next introduce an analogue to the definition of an ANR for fibrations that was originally defined by Yagasaki in [Yag85] and that can be seen as an ANR-type property for fiberwise spaces.
@@
The following result yields two helpful criteria to show the ANFR property for fibrations.
Lemma 5.5 ([Yag85, 1.2.(iii) and (v)]).
Let $Q$ and $B$ be metrizable spaces and let $f\colon Q\to B$ be a fibration.
  1. a)
    If $Q$ and $B$ are ANRs, then $f$ is an ANFR.
  2. b)
    If $f$ is a fiber bundle with typical fiber $F$ and if $F$ is an ANR, then $f\colon Q\to B$ is an ANFR.
@@
Put $D:=f^{-1}(C)$. Let $T$ be a metrizable space, let $A\subset T$ be closed and let $g\colon A\to D$ be continuous. Then $f\circ g\colon A\to B$ is continuous with $(f\circ g)(A)\subset C$. Since $C$ is an ANR, there exist an open neighborhood $V\subset T$ of $A$ and a continuous map $q\colon V\to B$ with $q|_{A}=f\circ g$ and $q(V)\subset C$. As an open subset of $T$, $V$ is again metrizable. Since $Q$ and $B$ are ANRs, $p$ is an ANFR by Lemma 5.5.a). Thus, there exist a neighborhood $U\subset V$ of $A$, which is open in $T$ and a continuous map $G\colon U\to Q$, such that
@@
In particular, it holds that $(f\circ G)(U)\subset q(U)\subset C$, which shows that $G(U)\subset D$. So we can restrict the codomain of $G$ to $G\colon U\to D$, which yields the desired continuous extension of $g$. Hence, $D$ is an ANR. ∎
@@
Since $E$ and $X$ are ANRs, $E^{r}$ and $X^{r}$ are again ANRs. Moreover, $p^{r}\colon E^{r}\to X^{r}$ is a fibration for which $E_{X}^{r}=(p^{r})^{-1}(\Delta_{r,X})$. Since $\Delta_{r,X}$ is homeomorphic to $X$, it is an ANR as well. As a special case of Lemma 5.6, we thus obtain that $E^{r}_{X}$ is an ANR.
Considering $E^{I}_{X}$, we first observe that since $p$ is a fibration the induced map of path spaces
$p_{\#}\colon PE\to PX,\quad p_{\#}(\gamma)=p\circ\gamma,$
is a fibration as well. Since $E$ and $X$ are ANRs, it further follows from [Hu65, Theorem VI.2.4] that $PE$ and $PX$ are ANRs. By definition of $E^{I}_{X}$, it holds that
$E^{I}_{X}=p_{\#}^{-1}(c(X)),$
where $c(X)\subset PX$ denotes the space of constant paths in $X$. Since $c(X)$ is homeomorphic to $X$, it is an ANR, so we again obtain from Lemma 5.6 that $E^{I}_{X}$ is an ANR. ∎
@@
By Proposition 5.7, $E^{r}_{X}$ is an ANR, so it follows from Lemma 5.5.a) that $q_{r}\colon E^{r}_{X}\to X$ is an ANFR. By Lemma 5.2, there exists a deformation $h^{\prime}\colon A\times[0,1]\to E^{r}_{X}$ with $h^{\prime}(a,1)=d_{E}(a)$ and $q_{r}(h^{\prime}(a,t))=q_{r}(a)$ for all $a\in A$ and $t\in[0,1]$, where $d_{E}(u_{1},u_{2},\dots,u_{r})=(u_{1},u_{1},\dots,u_{1})$. We put $Q^{\prime}:=(E^{r}_{X}\times\{0,1\})\cup(A\times[0,1])$ and let $H^{\prime}\colon Q^{\prime}\times[0,1]\to E^{r}_{X}$ be defined as in (2.2), only replacing $d$ by $d_{E}$ in the second case. Then $q_{r}\circ\mathrm{pr}_{1}|_{Q}=q_{r}\circ H^{\prime}$, where $\mathrm{pr}_{1}\colon E^{r}_{X}\times[0,1]\to E^{r}_{X}$ denotes the projection onto the first factor. Since $q_{r}$ is an ANFR, there exist an open neighborhood $V^{\prime}\subset E_{X}^{r}\times[0,1]$ of $Q^{\prime}$ and a continuous $K^{\prime}\colon V^{\prime}\to E_{X}^{r}$, such that $K^{\prime}|_{Q^{\prime}}=H$ and $q_{r}\circ\mathrm{pr}_{1}|_{V^{\prime}}=q_{r}\circ K^{\prime}$.
@@
    For all $A\subset B\subset E^{r}_{X}$, it holds that
    $\operatorname{\mathsf{TC}}_{r,p}(A)\leq\operatorname{\mathsf{TC}}_{r,p}(B).$
@@
    For all $A,B\subset E^{r}_{X}$ it holds that
    $\operatorname{\mathsf{TC}}_{r,p}(A\cup B)\leq\operatorname{\mathsf{TC}}_{r,p}(A)+\operatorname{\mathsf{TC}}_{r,p}(B).$
@@
All parts of the proposition follow in complete analogy with the corresponding parts of Proposition 2.6, where we use that by assumption, the fibers of $p$ are path-connected. Here, in the proof of d) we use that since $E$ is normal, $E^{r}$ is normal as well and so is its closed subspace $E^{r}_{X}$. ∎
@@
Then
$\operatorname{\mathsf{TC}}_{r,p}(\Phi_{1}(A))\geq\operatorname{\mathsf{TC}}_{r,p}(A).$
@@
Put $k:=\operatorname{\mathsf{TC}}_{r,p}(\Phi_{1}(A))$. In analogy with the proof of Proposition 2.7, one shows that there are closed subsets $B_{1},B_{2},\dots,B_{k}\subset\Phi_{1}(A)$ with $\bigcup_{j=1}^{k}B_{j}=\Phi_{1}(A)$, such that for each $j\in\{1,\ldots,k\}$ there exists a sequential parametrized motion planner $s_{j}\colon B_{j}\to PE$. Put $B_{j}^{\prime}:=\Phi_{1}^{-1}(B_{j})$ for each $j\in\{1,\ldots,k\}$, such that $A\subset\bigcup_{j=1}^{k}B_{j}^{\prime}$, and define a map $s_{j}^{\prime}\colon B_{j}^{\prime}\to PE$ by the same expression as in (2.3) in the proof of Proposition 2.7. One checks that condition (5.1) ensures that each $s_{j}^{\prime}$ is indeed a continuous sequential parametrized motion planner. Since $E$ and $X$ are ANRs, it follows from Lemma 5.8 that there are open subsets $U_{1},U_{2}\dots,U_{k}\subset E^{r}_{X}$ with $B^{\prime}_{j}\subset U_{j}$ for each $j\in\{1,2,\dots,k\}$, such that each $U_{j}$ admits a continuous sequential parametrized motion planner. Since $A\subset\bigcup_{j=1}^{k}U_{j}$, this shows that $\operatorname{\mathsf{TC}}_{r,p}(A)\leq k$. ∎

### v4

@@ 3. Navigation functions and sequential subspace complexities
For each $j\in\{1,2,\ldots,k\}$ we further put $\lambda_{j,1}:=1$. Choose a nowhere vanishing continuous vector field $X_{j}$ on $\mathbb{S}^{n_{j}}$. By rescaling we can arrange that $\|X_{j}(x)\|=(r-1)\cdot\pi$ for all $x\in\mathbb{S}^{n_{j}}$. With this normalization, $X_{j}(x_{j})$ is the tangent vector of a geodesic segment $a,a+\frac{1}{r-1}]\to\mathbb{S}^{n_{j}}$ from $x_{j}$ to $x_{j}$ for all $a\in\mathbb{R}$, $x_{j}\in\mathbb{S}^{n_{j}}$, $j\in\{1,2,\dots,k\}$. We define a sequential motion planner $s\colon\Sigma\to PM$ as follows:
@@
  • the constant path in $x_{j}$ if $\lambda_{j,\ell}=\lambda_{j,(\ell+1)}=-1$,
  • the unique length-minimizing geodesic segment from $x_{j}$ to $x_{j}$ with $\gamma^{\prime}(\frac{\ell-1}{r-1})=X_{j}(x_{j})$ if $\lambda_{j,\ell}=1$ and $\lambda_{j,(\ell+1)}=-1$,
  • the unique length-minimizing geodesic segment from $x_{j}$ to $x_{j}$ with $\gamma^{\prime}(\frac{\ell-1}{r-1})=X_{j}(-x_{j})$ if $\lambda_{j,\ell}=-1$ and $\lambda_{j,(\ell+1)}=1$.
One checks that the map $s$ we obtain in this way is indeed a continuous sequential motion planner on $\Sigma$. By Lemma 2.4 it extends to a continuous sequential motion planner on an open neighborhood of $\Sigma$. Thus, $\operatorname{\mathsf{TC}}_{r,M}(\Sigma)=1$ for all $\Sigma\in\mathcal{C}_{4i}(F)$ and $i\in\{0,1,\dots,k(r-1)\}$. Inserting this into (3.6) shows that $\operatorname{\mathsf{TC}}_{r}(\mathbb{S}^{n_{1}}\times\mathbb{S}^{n_{2}}\times\dots\times\mathbb{S}^{n_{k}})\leq k(r-1)+1$ for all odd numbers $n_{1},n_{2},\dots,n_{k}\in\mathbb{N}$.
@@ 4. Sequential topological complexity of sublevel sets
In this section, we want to derive upper bounds of the sequential TC of a sublevel set of a function $f\colon X\to\mathbb{R}$ from the results of section 2. More precisely, we will apply those results to functions $F\colon X^{r}\to\mathbb{R}$ of the form
@@
and will rephrase the upper bounds in terms of $f$. While we have previously studied subspace complexities of the form $\operatorname{\mathsf{TC}}_{r,X}(F^{\lambda})$, it turns out that we can find upper bounds of $\operatorname{\mathsf{TC}}_{r}(f^{\lambda})$, i.e. for the sequential TCs of $f^{\lambda}$ and not just for sequential subspace complexities.
@@
    It follows directly from the corresponding property of $(\varphi,f)$ that $(\Phi,F)$ satisfies ($D_{1}$), so we shall only prove that it satisfies ($D_{2}$). Let $B\subset X^{r}$ be such that $F$ is bounded on $B$ and $\inf\{F(x)-F(\Phi(x))\mid x\in B\}=0$. Let $(x_{n})_{n\in\mathbb{N}}$ be a sequence in $B$ with
@@
    Since $(F(x_{n}))_{n}$ is bounded and $f$ is bounded from below, it follows that $(f(x_{j,n}))_{n}$ is bounded for each $j$. By assumption, $(\varphi,f)$ satisfies ($D_{2}$), so $(x_{1,n})_{n}$ has a convergent subsequence $(x_{1,n_{k}})_{n\in\mathbb{N}}$. Likewise, $(x_{2,n_{k}})_{k}$ has a convergent subsequence $(x_{2,n_{k_{\ell}}})_{\ell}$. Going on like this until the $r$-th component sequence, we eventually obtain a subsequence of $(x_{n})_{n}$, which we simply denote by $(y_{m})_{m\in\mathbb{N}}$, all of whose component sequences converge. Thus, $(y_{m})_{m\in\mathbb{N}}$ is a convergent subsequence of $(x_{n})_{n\in\mathbb{N}}$ whose limit lies in $\overline{B}\cap\mathrm{Fix\;}\Phi_{1}$ by property ($D_{1}$). This shows that $(\Phi,F)$ has property ($D_{2}$).
@@
    By assumption the restriction $\Phi|_{A}\colon A\to A$ is well-defined. Since $(\Phi,F)$ satisfies condition ($D_{1}$) by a), it is evident that $(\Phi|_{A},F|_{A})$ satisfies ($D_{1}$). To show $(D_{2})$, let $B\subset A$ be a subset, such that $F$ is bounded on $B$ and
@@
    such that $\operatorname{cl}_{A}(B)\cap\mathrm{Fix\;}\Phi=\overline{B}\cap\mathrm{Fix\;}\Phi\neq\emptyset$, where $\operatorname{cl}_{A}(B)$ denotes the closure of $B$ in $A$.Hence, $(\Phi|_{A},F|_{A})$ satisfies $(D_{2})$.
@@
Apparently, since $f$ is bounded from below and since $\mathrm{Fix\;}\Phi=(\mathrm{Fix\;}\varphi)^{r}$, it holds for each $\mu\in\mathbb{R}$ that
$F_{\mu}\cap Y^{r}\cap\mathrm{Fix\;}\Phi$ $=\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,r\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}(f_{\mu_{i}}\cap Y\cap\mathrm{Fix\;}\varphi)=\bigcup_{\stackrel{{\scriptstyle\mu_{1},\dots,\mu_{r}\in(-\infty,\lambda]}}{{\mu_{1}+\dots+\mu_{r}=\mu}}}\prod_{i=1}^{r}(f_{\mu_{i}}\cap\mathrm{Fix\;}\Phi).$
@@ 5. Sequential parametrized subspace complexities as index functions
Throughout this section let $r\in\mathbb{N}$ with $r\geq 2$ and let $p\colon E\to X$ be a fibration with path-connected fibers. We further let $E^{I}_{X}$, $E^{r}_{X}$ and $\Pi_{r}$ be defined as in the introduction.
@@
    Let $A\subset E^{r}_{X}$. A local section $s\colon A\to E^{I}_{X}$ of $\Pi_{r}$ is called a sequential parametrized motion planner on $A$ .
@@
    The $r$-th sequential parametrized subspace complexity of $A$, denoted by $\operatorname{\mathsf{TC}}_{r,p}(A)$, is given as the minimal value of $n\in\mathbb{N}$ for which there exist open subsets $U_{1},U_{2},\dots,U_{n}\subset E^{r}_{X}$ with $A\subset\bigcup_{i=1}^{n}U_{i}$, such that there exists a continuous sequential parametrized motion planner $s_{i}\colon U_{i}\to E_{X}^{I}$ for each $i\in\{1,2,\dots,n\}$. In case there is no such $n$, we put $\operatorname{\mathsf{TC}}_{r,X}(A):=\infty$ .
Remark 5.2.
  1. (1)
    Let $B\subset E$ and consider its $r$-fold fiber product $B_{X}^{r}\subset E^{r}_{X}$. Assume that the restriction $p|_{B}\colon B\to X$ is again a fibration. Similarly to the unparametrized case, see Remark 2.2.(1), we note that one needs to distinguish between the $r$-th sequential parametrized subspace complexity $\operatorname{\mathsf{TC}}_{r,p}(B_{X}^{r})$ and the value of $\operatorname{\mathsf{TC}}_{r}[p|_{B}\colon B\to X]$.
    As an example, let $X=M$ be a closed manifold and consider its tangent bundle $p\colon TM\to M$. Since the fibers of $p$ are vector spaces, one obtains that $\operatorname{\mathsf{TC}}_{r}[p\colon TM\to M]=1$, see [CFW21, Proposition 4.5]. Choose a Riemannian metric $g$ on $M$ and let $UM$ be its unit tangent bundle. The restriction $p|_{UM}\colon UM\to M$ is a fibration as well. Restricting the global sequential parametrized motion planner of the tangent bundle shows that $\operatorname{\mathsf{TC}}_{r,p}(UM_{M}^{r})=1$. However, $\operatorname{\mathsf{TC}}_{r}[p|_{UM}\colon UM\to M]\geq\operatorname{\mathsf{TC}}_{r}(\mathbb{S}^{\mathrm{dim}(M)-1})$ since its fiber is the sphere $\mathbb{S}^{\mathrm{dim}(M)-1}$, see [FP22]. It is well-known that $\operatorname{\mathsf{TC}}_{r}(\mathbb{S}^{\mathrm{dim}(M)-1})\geq 2$, so the sequential parametrized subspace complexity differs from the intrinsic sequential parametrized topological complexity of $p|_{B}$.
  2. (2)
    As elaborated upon in the introductions of [CFW21] and [FP22], sequential parametrized topological complexity is motivated by robot motion planning inside a workspace which depends on certain parameters, e.g. the positions of certain obstacles on a warehouse floor. Here, the selection of a parameter corresponds to the selection of a fiber of $p\colon E\to X$, in which motion planning algorithms are carried out. In complete analogy with sequential subspace complexities, see Remark 2.2.(2), sequential parametrized subspace complexities correspond to the selection of a subset of $E^{r}_{X}$, i.e. a certain set of $r$-tuples of points, such that all points in a tuple lie in the same fiber of $p$, instead of considering motion planning algorithms covering all of $E^{r}_{X}$.
@@
More abstractly, $q_{r}$ is given as the pullback fibration $q_{r}=\Delta_{r}^{*}p^{r}$, where $\Delta_{r}\colon X\to X^{r}$, $\Delta(x)=(x,x,\dots,x)$ and where $p^{r}\colon E^{r}\to X^{r}$ denotes the $r$-fold cartesian product of $p$ with itself. In particular, $q_{r}$ is itself a fibration.
@@
    AA admits a continuous sequential parametrized motion planner.
@@
so $s(a)\in E^{I}_{X}$ for all $a\in A$. Thus, $s$ is a continuous sequential parametrized motion planner. ∎
In section 2 we have restricted ourselves to considering ANRs when considering sequential subspace complexities, as we could establish the continuity property of an index function under this assumption. Likewise, we need to make a similar assumption on $p$ to ensure that sequential parametrized subspace complexities will have the continuity property. For this purpose, we next introduce an analogue to the definition of an ANR for fibrations that was originally defined by Yagasaki in [Yag85] and that can be seen as an ANR-type property for fiberwise spaces.
@@
The following result yields a helpful criterion to show the ANFR property for fibrations.
Lemma 5.6 ([Yag85, 1.2.(iii)]).
Let $Q$ and $B$ be topologial spaces and let $f\colon Q\to B$ be a fibration. If $Q$ and $B$ are ANRs, then $f$ is an ANFR.
@@
Put $D:=f^{-1}(C)$. Let $T$ be a metrizable space, let $A\subset T$ be closed and let $g\colon A\to D$ be continuous. Since $(f\circ g)(A)\subset C$ and $C$ is an ANR, there exist an open neighborhood $V\subset T$ of $A$ and a continuous map $q\colon V\to B$ with $q|_{A}=f\circ g$ and $q(V)\subset C$. As an open subset of $T$, $V$ is again metrizable. Since $Q$ and $B$ are ANRs, $p$ is an ANFR by Lemma 5.6. Thus, there exist a neighborhood $U\subset V$ of $A$, which is open in $T$ and a continuous map $G\colon U\to Q$, such that
@@
Hence, $(f\circ G)(U)\subset q(U)\subset C$, which shows that $G(U)\subset D$. So we can restrict the codomain of $G$ to $G\colon U\to D$, which yields a continuous extension of $g$. Hence, $D$ is an ANR. ∎
@@
Since $E$ and $X$ are ANRs, $E^{r}$ and $X^{r}$ are again ANRs. Moreover, $p^{r}\colon E^{r}\to X^{r}$ is a fibration for which $E_{X}^{r}=(p^{r})^{-1}(\Delta_{r,X})$. Since $\Delta_{r,X}$ is homeomorphic to $X$, it is an ANR as well. As a special case of Lemma 5.7, we thus obtain that $E^{r}_{X}$ is an ANR. Considering $E^{I}_{X}$, we observe that the induced map of path spaces $p_{\#}\colon PE\to PX$, $p_{\#}(\gamma)=p\circ\gamma$, is a fibration as well and it follows from [Hu65, Theorem VI.2.4] that $PE$ and $PX$ are ANRs. By definition of $E^{I}_{X}$, it holds that $E^{I}_{X}=p_{\#}^{-1}(c(X))$, where $c(X)\subset PX$ denotes the space of constant paths in $X$. Since $c(X)$ is homeomorphic to $X$, it is an ANR, so we obtain from Lemma 5.7 that $E^{I}_{X}$ is an ANR. ∎
@@
By Proposition 5.8, $E^{r}_{X}$ is an ANR, so it follows from Lemma 5.6 that $q_{r}\colon E^{r}_{X}\to X$ is an ANFR. By Lemma 5.3, there exists a deformation $h^{\prime}\colon A\times[0,1]\to E^{r}_{X}$ with $h^{\prime}(a,1)=d_{E}(a)$ and $q_{r}(h^{\prime}(a,t))=q_{r}(a)$ for all $a\in A$ and $t\in[0,1]$, where $d_{E}(u_{1},u_{2},\dots,u_{r})=(u_{1},u_{1},\dots,u_{1})$. We put $Q^{\prime}:=(E^{r}_{X}\times\{0,1\})\cup(A\times[0,1])$ and let $H^{\prime}\colon Q^{\prime}\times[0,1]\to E^{r}_{X}$ be defined as in (2.2), only replacing $d$ by $d_{E}$ in the second case. Then $q_{r}\circ\mathrm{pr}_{1}|_{Q}=q_{r}\circ H^{\prime}$, where $\mathrm{pr}_{1}\colon E^{r}_{X}\times[0,1]\to E^{r}_{X}$ denotes the projection onto the first factor. Since $q_{r}$ is an ANFR, there exist an open neighborhood $V^{\prime}\subset E_{X}^{r}\times[0,1]$ of $Q^{\prime}$ and a continuous $K^{\prime}\colon V^{\prime}\to E_{X}^{r}$, such that $K^{\prime}|_{Q^{\prime}}=H$ and $q_{r}\circ\mathrm{pr}_{1}|_{V^{\prime}}=q_{r}\circ K^{\prime}$.
@@
    For all $A\subset B\subset E^{r}_{X}$, it holds that $\operatorname{\mathsf{TC}}_{r,p}(A)\leq\operatorname{\mathsf{TC}}_{r,p}(B)$.
@@
    For all $A,B\subset E^{r}_{X}$ it holds that $\operatorname{\mathsf{TC}}_{r,p}(A\cup B)\leq\operatorname{\mathsf{TC}}_{r,p}(A)+\operatorname{\mathsf{TC}}_{r,p}(B)$.
@@
All parts of the proposition follow in complete analogy with the corresponding parts of Proposition 2.5, where we use that by assumption, the fibers of $p$ are path-connected. Here, in d) we use that since $E$ is normal, $E^{r}$ is normal as well and so is its closed subspace $E^{r}_{X}$. ∎
@@
Then $\operatorname{\mathsf{TC}}_{r,p}(\Phi_{1}(A))\geq\operatorname{\mathsf{TC}}_{r,p}(A)$.
@@
This follows along the lines of the proof of Proposition 2.6, using Lemma 5.9 instead of Lemma 2.4. Note that condition (5.1) ensure that the maps $s_{j}^{\prime}$ that are given as in (2.3), are indeed sequential parametrized motion planners. ∎